In a game, a card is taken at random from a full pack of 52 playing cards. It is then replaced, and a second card is taken. What is the probability that both cards are diamonds? O 0.225 O 0.625 O 0.0625 O 0.5 O 0.125


1
Expert's answer
2021-09-04T01:01:06-0400

Given information-

We have given a deck of 52 playing cards.

We know that we have 13 diamonds cards in a deck of 52 playing cards.

One by one we have to draw a card with a replacement.

We have to find the probability that both cards are diamonds.


Let A be the event that selecting diamonds cards.

P(Selecting a diamond card at first draw) = P(A) = 13/52

 

Since we replace a card so the pack still remains 52 cards and, then we take a second card from the pack.

P(Selecting a diamond card at second draw) = P(B) = 13/52

 

So the probability that drawing a diamond card for both the first and second draw is (13/52)*(13/52)

P(Drawing both cards are diamonds) = P(A)*P(B)= (13/52)*(13/52) = 1/16= 0.0625.

P(Drawing both cards are diamonds) = 0.0625.

Option C is the correct answer.